On January 1, Revis Consulting entered into a contract to complete a cost reduction program for Green Financial over a six-month period. Revis will receive $22,400 from Green at the end of each month. If total cost savings reach a specific target, Revis will receive an additional $11,200 from Green at the end of the contract, but if total cost savings fall short, Revis will refund $11,200 to Green. Revis estimates an 80% chance that cost savings will reach the target and calculates the contract price based on the expected value of future payments to be received. Required: Prepare the following journal entries for Revis: 1. to 3. Prepare the journal entry on January 31 to record the collection of cash and recognition of the first month’s revenue, assuming total cost savings exceed target, record the entry on June 30 for receipt of the bonus and assuming total cost savings fall short of target and record the entry on June 30 for payment of the penalty. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

Journal entries:

1.

Date General Journal Debit Credit
Jan. 31 Cash $22,400
Bonus receivable ($23,520 - 22,400) 1,120
Service revenue ($141,120 / 6 ) $23,520
(To record collection of cash and revenue recognition)

Working note:

$(22,400 * 6) + $11,200 * 80% = $116,480

$(22,400 * 6) - $11,200 * 20% = $24,640

Total expected value = $116,480 + 24,640 = $141,120

2.

Date General Journal Debit Credit
June 30 Cash $11,200
Bonus receivable ($1,120 * 6) $6,720
Service revenue ($11,200 - 6,720) $4,480
(To record receipt of bonus)

3.

Date General Journal Debit Credit
June 30 Service revenue $17,920
Bonus receivable 6,720
Cash $11,200
(To record payment of penalty)